OpenStudy (anonymous):

QUESTION 1 1. If f(x) = x2 and g(x) = x + 3, find (f + g)(x). Path: pWords:0 3 points QUESTION 2 2. If f(x) = x2 and g(x) = x + 3, find (f - g)(x). Path: pWords:0 3 points QUESTION 3 3. If f(x) = x2 and g(x) = x + 3, find (fg)(x). Path: pWords:0 3 points QUESTION 4 4. If f(x) = x - 1 and g(x) = 2x - 3, find (f + g)(x). Path: pWords:0 3 points QUESTION 5 5. If f(x) = x - 1 and g(x) = 2x - 3, find (f - g)(x). Path: pWords:0 3 points QU

OpenStudy (mayankdevnani):

is your question look like this :- 1) f(x)+g(x)

OpenStudy (anonymous):

for which

OpenStudy (mayankdevnani):

for your question 1

OpenStudy (anonymous):

f(x) = x2 and g(x) = x + 3, find (f + g)(x).

OpenStudy (mayankdevnani):

(f+g)(x) doesn't make sense rather it would be like this f(x)+g(x)

OpenStudy (mayankdevnani):

do you know the answer ? i can confirm it then

OpenStudy (anonymous):

I'm trying to solve it I'm having trouble

OpenStudy (mayankdevnani):

do you know the answer ?

OpenStudy (anonymous):

no

OpenStudy (daniellelovee):

just add the equations and subtract them :)

OpenStudy (anonymous):

could you show me one? just to be sure i got this

OpenStudy (mayankdevnani):

ok! i'll show you

OpenStudy (mayankdevnani):

for question 1 \[\large \bf f(x)=x^2~~~and~~~g(x)=x+3\] \[\large \bf (f+g)x=f(x)+g(x)\] All you need to do is just add both equations.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

and then?

OpenStudy (mayankdevnani):

so what you get ?

OpenStudy (mayankdevnani):

you need to add them only,you'll get your answer

OpenStudy (mayankdevnani):

so tell me what you get after adding both equations ?

OpenStudy (anonymous):

do you just add the

OpenStudy (mayankdevnani):

okay i'll do one for you

OpenStudy (mayankdevnani):

f(x)+g(x)=x^2+x+3=ANSWER

OpenStudy (mayankdevnani):

understood ?

OpenStudy (anonymous):

yes

OpenStudy (mayankdevnani):

great

OpenStudy (mayankdevnani):

now move on to second one

OpenStudy (anonymous):

so your answer is one number?

OpenStudy (mayankdevnani):

nope

OpenStudy (mayankdevnani):

answer is whole term

OpenStudy (mayankdevnani):

\[\large \bf \boxed{x^2+x+3}=\color{red}{Answer}\]

OpenStudy (anonymous):

ok thanks

OpenStudy (mayankdevnani):

now can we move on to second question ? can we ?

OpenStudy (anonymous):

yes

OpenStudy (mayankdevnani):

\[\large \bf (f-g)(x)=f(x)-g(x)\] all you need to do is subtract the equations

OpenStudy (mayankdevnani):

just plug in the values of f(x) and g(x) in above equation,you'll get your answer

OpenStudy (anonymous):

f x 2- 3?

OpenStudy (mayankdevnani):

kinda like that

OpenStudy (mayankdevnani):

so your answer to second question is :- x^2-(x+3)=x^2-x-3

OpenStudy (anonymous):

I'm sorry I'm really bad at these

OpenStudy (mayankdevnani):

lol no problem

OpenStudy (anonymous):

so i missed the x

OpenStudy (mayankdevnani):

yep

OpenStudy (anonymous):

thanks